Two kids are playing tag one kid is travelling at a constant velocity of 0.5m/s. the other kid starts from rest and accelerates at 1.2m/s2. How long does it take the other kid to catch up.
1
Expert's answer
2018-10-23T11:25:09-0400

s_1=vt

s_2=(a·t^2)/2

If s_1=s_2 vt=(a·t^2)/2

t=2v/a=(2·0.5)/1.2=0.833 sec

s_1=s_2=0.5·0.833=0.4167 metres